OpenStudy (anonymous):

Please help me out. :) lim (2-3y^2) / (5y^2+4y) x approaches positive infinity

ganeshie8 (ganeshie8):

divide top and bottom by y^2

ganeshie8 (ganeshie8):

\(\large \lim \limits_{y\to \infty} ~ \dfrac{2-3y^2}{5y^2+4y}\)

OpenStudy (anonymous):